1. Suppose Mary intends to sell two software products X & Y for the next convention &



Units Sold.6040100



Required:What is the BEP (in units & in Birr



Fixed Costs4,500



Operating Income$ 1,500



budgets the following.



Variable Costs,$120$70 per unit7,2002,80010,000



1.What is the sales mix of videos and equipment sets



Unit Contribution Margin,$80$ 30per unit$ 4,800$ 1200$ 6,000



Revenues,$200$100 per unit$12,000$ 4,000$16,000



Required: Answer the following



XYTotal



2.Compute weighted average contribution margin

1
Expert's answer
2022-03-21T15:52:03-0400
  1. $4500
  2. (80+30+4800+1200+600)/5 = 1342
  3. 70-30= 40USD
  4. 80+30+4800+1200+600)/5 = 1342
  5. 4500-1500 = 3000.